[figure: 9.7 cm]

STUDY OF HIGH - T_{C} SUPERCONDUCTOR WITH FINE MAGNETIC PARTICLES COMPOSITES

M. Timko^{a}, P. Kopčanský^{a}, J. Kovač^{a}, M. Koneracká^{a}, V. Plecháček ^{b}

aInstitute of Experimental Physics, Slovak Academy of Sciences, Watsonova 47, 043 53 Košice, Slovakia bInstitute of Physics, Czech Academy of Sciences, Cukrovarnícka 10, CZ - 162 00 Praha 6, Czech Republic

The magnetic state of magnetic material in a composite system of magnetic and high T_{c} - superconducting powder affects the flux exclusion properties of superconducting materials. The composite systems as a mixture of fine magnetic particles of F e_{3}O_{4} (10 nm in diameter) and superconducting powder (particle size 5 - 15μm) such as Bi_{1.8}P b_{0.2}Sr_{2}Ca_{2}Cu_{3}O_{10+x} and Y Ba_{2}Cu_{3}O_{7−x} have been prepared. The influence of internal magnetic field (generated by magnetic particles) on the superconducting properties of prepared composites have been studied as a function of concentration of magnetic particles and their magnetic state. The zero field and field cooling magnetization as a function of temperature and magnetic field was measured using standard VSM magnetometer. The magnetic flux exclusion properties of superconducting material were affected by the presence of internal magnetic field larger than the lower critical field of superconducting material. The experimental data were discussed in the framework of the theory where a distribution of internal magnetic field was taken into account.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : M. Timko

Address for correspondence : Institute of Experimental Physics, SAS Watsonova 47, 043 53 Košice Slovakia

Email address : timko@saske.sk